OpenStudy (anonymous):

a car starting from rest moves with a constant acceleration of 2.0m/s^2 for 10s, how far does it travel

OpenStudy (anonymous):

s=0.5 *a *t^2...u have a and t ,..find s.

OpenStudy (anonymous):

s=ut+1/2a t^2 u=0 s=1/2*2*10*10 =100m